# Larger Scale Offers Better Security in the Nakamoto-style Blockchain

### v3

@@ Abstract.
Traditional security models for Nakamoto-style blockchains assume instantaneous synchronization among malicious nodes, which overestimate adversarial coordination capability. We revisit these existing models and propose two more realistic security models. First, we propose the static delay model. This model first incorporates adversarial communication delay. It quantifies how the delay constrains the effective growth rate of private chains and yields a closed-form expression for the security threshold. Second, we propose the dynamic delay model that further captures the decay of adversarial corruption capability and the total adversarial delay window. Theoretical analysis shows that private attacks remain optimal under both models. Finally, we prove that large-scale Nakamoto-style blockchains offer better security. This result provided a theoretical foundation for optimizing consensus protocols and assessing the robustness of large-scale blockchains.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: the 2026 ACM SIGSAC Conference on Computer and Communications Security; November 15-19, 2026; The Hague, The Netherlandsisbn: 978-1-4503-XXXX-X/2026/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1.4. Proposed Solutions
Based on this derivation, the model reveals the critical relationship between the communication delay $\Delta_{a}$ of adversarial nodes and the security threshold $\beta^{*}$ (the maximum proportion of adversarial node computing power that the system can tolerate): an increase in $\Delta_{a}$ can significantly raise the security threshold $\beta^{*}$. When the communication delay $\Delta_{a}$ of adversarial nodes is greater than the communication delay $\Delta$ of honest nodes, $\beta^{*}$ may even exceed the boundary of the “51% attack” in traditional blockchain security models (i.e., $\beta^{*}>0.5$). This implies that in scenarios where the synchronization efficiency of adversarial nodes is lower, the system may still remain secure even if adversarial nodes control more than 50% of the computing power.
@@ 1. $\bullet$
    Delay-Driven Security Threshold Correction. The static $\Delta,\Delta_{a}$ model quantifies how $\Delta_{a}$ increases $\beta^{*}$, breaking the 51% boundary when $\Delta_{a}>\Delta$. We derive $\beta^{*}\approx\frac{1}{2}+\frac{\lambda(\Delta_{a}-\Delta)}{4}$, establishing a monotonic relationship between $\beta^{*}$ and $\Delta_{a}$.
@@ 2. $\bullet$
    Security of Large-Scale Blockchains. The dynamic model demonstrates that adversarial power decay ensures asymptotic security superiority for large networks, with $\Pr[\beta\leq\beta^{*}]\rightarrow 1$ as $n\rightarrow\infty$.
@@ 3. $\bullet$
    Optimality of Private Attacks. We formally prove that private attacks remain optimal in both models.
@@ Appendix I Proof of Theorem 4
This approximation aligns with the ”power dilution” principle in decentralized systems, where $\beta$ becomes increasingly predictable as $n$ grows.
@@
The interplay among logarithmic threshold decay ($\beta^{*}$), polynomial corruption suppression ($p^{*}(n)$), and network scaling ($\sqrt{n}$) creates a multiplicative effect. This aligns with the ”security through scale” axiom in decentralized systems, where larger networks amplify the adversary’s coordination costs while diluting their power.

### v4

@@ Abstract.
Security analyses of Nakamoto-style blockchains typically assume that adversarial nodes can coordinate instantaneously through private channels. This assumption substantially overestimates adversarial effectiveness in real-world deployments, where adversaries are geographically distributed and subject to non-negligible coordination and communication delays.
In this paper, we revisit the security of Nakamoto-style consensus under explicit adversarial coordination constraints. We first introduce a static delay model that treats adversarial internal communication latency as an independent parameter. Using queueing theoretic analysis, we derive the effective growth rate of adversarial private chains and obtain a corrected security threshold. The results show that adversarial coordination delay monotonically improves security and can exceed the classical 50% bound when adversarial coordination is slower than honest block propagation. We then develop a dynamic delay corruption model that captures the joint efforts of network scale, communication delay, and probabilistic node corruption under bounded adversarial resources. Under this model, we prove that increasing network scale simultaneously amplifies adversarial coordination costs and dilutes adversarial power, leading to asymptotically improved security: the probability that adversarial power exceeds the security threshold converges to zero as the network grows. Finally, we show that private chain attacks remain optimal whthin the considered threat model. Coordination delay constrains adversarial effectiveness without introducing more powerful attacks. Our findings provide a theoretical foundation for optimizing consensus protocols and assessing the robustness of large-scale blockchains.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: the 2026 ACM SIGSAC Conference on Computer and Communications Security; November 15-19, 2026; The Hague, The Netherlandsisbn: 978-1-4503-XXXX-X/2026/06
@@ 1.4. Proposed Solutions
Based on this derivation, the model reveals the critical relationship between the communication delay $\Delta_{a}$ of adversarial nodes and the security threshold $\beta^{*}$ (the maximum proportion of adversarial node computing power that the system can tolerate): an increase in $\Delta_{a}$ can significantly raise the security threshold $\beta^{*}$. When the communication delay $\Delta_{a}$ of adversarial nodes is greater than the communication delay $\Delta$ of honest nodes, $\beta^{*}$ may even exceed the boundary of the 51% attack in traditional blockchain security models (i.e., $\beta^{*}>0.5$). This implies that in scenarios where the synchronization efficiency of adversarial nodes is lower, the system may still remain secure even if adversarial nodes control more than 50% of the computing power.
@@ 1. $\bullet$
    Modeling adversarial coordination delay. We introduce adversarial internal communication delay as a first-class parameter in Nakamoto-style security analysis, separating it from honest network delay and capturing coordination inefficiencies inherent in geographically distributed adversaries.
@@ 2. $\bullet$
    Static delay model and corrected security threshold. We propose a static delay model incorporating adversarial coordination latency and derive the effective growth rate of adversarial private chains using queueing-theoretic analysis. This yields a closed-form approximation of the security threshold.
@@ 3. $\bullet$
    Dynamic delay–corruption model for scaling networks. We develop a dynamic model that captures the joint effects of network scale, communication delay, and probabilistic node corruption under bounded adversarial resources. The model formalizes how scaling simultaneously increases adversarial coordination cost and dilutes adversarial power.
  4. $\bullet$
    Asymptotic security characterization. exceeding the security threshold converges to zero as the network size grows. This result provides a formal explanation for how decentralization can enhance security under realistic coordination constraints.
  5. $\bullet$
    Optimality of private attacks under coordination delay. We show that private attacks remain optimal adversarial strategies within the considered threat model, establishing that coordination delay constrains adversarial effectiveness without introducing more powerful attack classes.
@@ Appendix I Proof of Theorem 4
This approximation aligns with the power dilution principle in decentralized systems, where $\beta$ becomes increasingly predictable as $n$ grows.
@@
The interplay among logarithmic threshold decay ($\beta^{*}$), polynomial corruption suppression ($p^{*}(n)$), and network scaling ($\sqrt{n}$) creates a multiplicative effect. This aligns with the “security through scale” axiom in decentralized systems, where larger networks amplify the adversary’s coordination costs while diluting their power.